OpenStudy (cedric):

After a discount of 10%, the price of a jacket, including tax, is $70.79. The following equation can be used to find x, the price of the jacket without the discount. x - 0.1x + 0.06(x - 0.1x) = 70.79 What is the price of the jacket without the discount?

OpenStudy (anonymous):

70.79

OpenStudy (ganpat):

x - 0.1x + 0.06(x - 0.1x) = 70.79 .. can u isolate x ?? take it one side and rest all on other..

OpenStudy (cedric):

ok hold on

OpenStudy (anonymous):

It is not about the variable X, ok.

OpenStudy (ganpat):

@Zekarias : After a discount of 10%, the price of a jacket, including tax, is $70.79... so price of jacket after discount is 70.79$

OpenStudy (cedric):

so it'd be 74.20?

OpenStudy (anonymous):

Thank you you are correct, cedric.

OpenStudy (ganpat):

x - 0.1x + 0.06(x - 0.1x) = 70.79 x - 0.1x + 0.06x - 0.006x = 70.79 so x = 74.20.. $.. u r right..

OpenStudy (cedric):

alright thanks you guys